# Stylized synthetic augmentation further improves Corruption Robustness

### v2

@@ Abstract
This paper proposes a training data augmentation pipeline that combines synthetic image data with neural style transfer in order to address the vulnerability of deep vision models to common corruptions. We show that although applying style transfer on synthetic images degrades their quality with respect to the common FID metric, these images are surprisingly beneficial for model training. We conduct a systematic empirical analysis of the effects of both augmentations and their key hyperparameters on the performance of image classifiers. Our results demonstrate that stylization and synthetic data complement each other well and can be combined with popular rule-based data augmentation techniques such as TrivialAugment, while not working with others. Our method achieves state-of-the-art corruption robustness on several small-scale image classification benchmarks, reaching 93.54%, 74.9% and 50.86% robust accuracy on CIFAR-10-C, CIFAR-100-C and TinyImageNet-C, respectively.
@@ 1 INTRODUCTION
In recent years, computer vision models have surpassed human performance in tasks such as image classification. However, unlike the human vision system, such models tend to be unstable when faced with small changes to their input data. This is true for both adversarially optimized perturbations [goodfellow2014] and randomly distributed real-world image corruptions [Hendrycks2019a]. Achieving robustness to such changes is an essential technical property of vision models and even a regulatory prerequisite for safety critical applications [EU_AI_Act_2024].
@@
This study examines the combination of two model-guided data augmentation approaches, synthetic images and Neural Style Transfer (NST), as illustrated in Figure 1. Synthetic images from generative models still exhibit an appearance gap to original images, limiting their effective use in model training [wang2024improving]. We hypothesize that NST can circumvent this gap by exchanging potentially flawed synthetic image textures, making a model learn corruption robust features from high-quality synthetic object shapes instead.
@@
    We show that NST and synthetic images combine well: Although NST degrades the quality of synthetic images according to the popular FID metric, stylizing synthetic images improves the model performance and is empirically more effective than stylizing original images.
@@ 2 BACKGROUND
where $x,y$ are sampled from a data distribution $D$ and $C$ is a set of corruptions that alter $x$. $C$ can contain real-world [Hendrycks2019a], particularly challenging [Mintun2021], or theoretically motivated image corruptions [wang2021statistically, Siedel2022]. A straight-forward way to measure corruption robustness in (1), is to average test accuracies over all corruption sets.
@@
FID: Let $X$ be a random vector in a finite-dimensional vector space. We denote its mean as $\mu_{X}=\mathbb{E}[X]$ and its variance as $C_{X}=\mathbb{E}[(X-\mu_{X})(X-\mu_{X})^{\top}]$. Given two random vectors $X$ and $X^{\prime}$, the Fréchet Inception Distance (FID) given by
@@ 3.1 Rule-based data augmentation
Mixup [Zhang2018mixup] and Cutmix [Yun2019cutmix] interpolate between two images and their respective labels for training. Mixup boosts model robustness by enforcing a smooth, linear behavior. Noisy Feature Mixup (NFM) extends Mixup and Noise Injections into the feature space, improving their positive effect. NFM is only one example where combining rule-based methods yields superior results. Pytorch’s top image classification checkpoints use TrivialAugment, Cutmix, Mixup, and Random Erasing [Vryniotis2021]. NoisyMix [Erichson2024] is the top ranking method for achieving corruption robustness for the CIFAR-10/100 datasets on Robustbench [croce2020b] and combines NFM with AugMix and the Jensen–Shannon Divergence (JSD) loss that enforces consistency between augmented samples.
@@ 3.2.1 Synthetic Data Augmentation
Synthetic training data has been applied for both large-scale training [he2023] an in data-scarce fields such as medicine [Frid2018] to boost model performance. Generative models can therefore assist training on downstream tasks such as image classification [azizi2023]. For classifier training, class-conditional image generation, e.g. with [dhariwal2021] or without [ho2021] guidance of a classifier or via text prompts [ramesh2022] enables labeled data generation, thereby eliminating the need for pseudo-labeling [gowal2020uncovering].
Generative adversarial networks (GANs) have long been used for synthetic data augmentation [zheng2017unlabeled]. Later, diffusion models [ho2020denoising] surpassed GANs by delivering superior image quality [rombach2022, karras2022elucidating], typically measured through a lower FID score (see section 2). However, there still exists a significant domain gap between synthetic and original data due to a discrepancy in quality and appearance of synthetic images [wang2024improving]. This gap limits the effectiveness of synthetic data for training.
Several studies focus on improving the adversarial robustness of classifiers in particular through the addition of synthetic data during training [gowal2021improving, rebuffi2021fixing, wang2023better]. In fact, all top adversarially robust models on the popular Robustbench leaderboard [croce2020b] leverage generated data. Although some works hint at benefits for domain- and out-of-distribution generalization [Sandfort2019, wu2023datasetdm], the effect of additional generated training data on corruption robustness as defined in section 2 has to the best of our knowledge not been studied extensively. The top corruption robust models on Robustbench rely solely on rule-based data augmentation [croce2020b].
@@ 3.2.2 Neural Style Transfer
NST was used as a data augmentation method by [geirhos2018imagenettrained] when introducing a stylized version of the ImageNet dataset. Training with this dataset improves corruption robustness by forcing a convolutional neural network (CNN) to recognize object shapes instead of overfitting on textures, as those are random after stylization. However, [geirhos2018imagenettrained] trained on a fixed stylized version of the dataset. Ideally, we want an adaptive data augmentation scheme that allows a new stylization view every training epoch. [jackson2019style] did such data augmentation using the NST approach in [ghiasi2017exploring]. However, they did not evaluate corruption robustness. Neither of the two works have used their methods to fix the appearance gap of synthetic training data by exchanging unrealistic textures.
@@ 4 METHODOLOGY
Our data augmentation approach is to the best of our knowledge the first to explore a combination of synthetic image data augmentation and NST for maximizing the corruption robustness of image classfiers. To this end, our approach makes use of original, synthetic and style images as can be seen from Figure 1.
@@ 4.2 Stylization
  • $\lambda_{o}$ and $\lambda_{s}$ are independent probabilities that a synthetic image or an original image are stylized, respectively (see Figure 1).
@@
Table 1: FID of synthetic images over stylization ratio $\lambda_{s}$ for 50.000 synthetic images (100.000 for TIN). $\alpha_{s}$ is random uniformly drawn from [0.1,1.0] as is optimal in our experiments. Original denotes the FID before any pre-selection against the original reference.
@@ 5.2 Stylization ruins FID but improves model training
Let $D$ be the marginal feature distribution of the original training data. The mean and variance of the features $X$ of the original training data is given by $\mu=\mathbb{E}_{X\sim D}[X]$ and $C=\mathbb{E}_{X\sim D}[(X-\mu)(X-\mu)^{\top}]$. The mean and variance of the feature distribution $D^{\prime}$ of the synthetic dataset is given by $\mu^{\prime}$ and $C^{\prime}$. Let $D^{\prime}_{s}$ be a joint distribution of style images and synthetic data.
@@
Figure 2: Test accuracy, robustness and FID of stylized synthetic images plotted over varying synthetic stylization ratios $\lambda_{s}$. The accuracy and robustness maxima are not aligned with the minimum FID of the synthetic training data. Hence, FID is not a good predictor for the usefulness of synthetic images for data augmentation.
@@ 5.3 Parameter Tuning
Parameter $\alpha$: Figure 3 plots validation‑set tuning of $\alpha_{o}$ and $\alpha_{s}$ for TinyImageNet. For all our dataset, contrary to results in [jackson2019style] where $\alpha=0.5$ is optimal, we achieve best accuracy and robustness on original data at $\alpha_{o}=1.0$. On synthetic data, $\alpha_{s}$ is in a trade‑off, where increasing it too far boosts robustness but reduces accuracy. We choose to select $\alpha_{s}$ randomly drawn from $[0.1,1.0]$ as a balanced choice.
@@
Figure 5: Robustness and accuracy maps depending on the ratio of stylization on original ($\lambda_{o}$, x-axis) and synthetic ($\lambda_{s}$, y-axis) images. Values are averaged across 5 runs, synthetic image ratio $\lambda$ is 0.5. The points mark the lambda combination that is optimal for that dataset with respect to the mean of accuracy and robustness. On CIFAR in general, but also for optimal robustness on TIN, $\lambda_{o}<\lambda_{s}$.
@@
Population based training:
@@
In addition to the fragmentary grid search for individual hyperparameters, we generate an optimized augmentation schedule using population based training (PBT) [jaderberg2017population, ho2019population]. PBT trains multiple trials in parallel, promotes successful over unsuccessful checkpoints every predefined interval and perturbs the augmentation hyperparameters. It then tracks the most effective schedule over the training epochs. Figure 7 shows the resulting schedules for the lambda parameters, while the alpha parameters are set to constant values as derived from prior tuning. We apply PBT on train-validation data without resampling of parameters, with a perturbation interval of 1 epoch, after which trials are ranked according to the mean of accuracy and robustness, and 24 overall trials. We then replayed the schedule on train-test data. The schedule confirms for the CIFAR datasets that in tendency, synthetic data can take more stylization then original data, hence $\lambda_{s}>\lambda_{o}$. Towards the low learning rate sections of the training schedule at 140 and 300 epochs, there is a tendency towards more policy changes and higher stylization probabilities. The best synthetic ratio $\lambda$ clearly trends down from the start value of 0.5 that was found optimal before.
Table 2: Ours (SYN+NST+TA) outperforms other data augmentation methods. *NoisyMix is a 300 epoch rerun of the original implementation without early stopping, IPMix are original paper results. Both feature different parameters, IPMix uses the deeper WRN-40-4 architecture [Erichson2024, huang2024ipmix].
@@ 5.4 Ablation and Comparison Studies
Table 4 shows further ablations, adding popular rule‑based augmentations to the SYN+NST+TA baseline. Adding Random Erasing and noise in input and feature space in most cases further improves robustness. Prior work demonstrated that the listed additional approaches synergize well in a purely rule‑based setting [Vryniotis2021, Erichson2024]. However, combined with our approach, Mixup and CutMix degrade performance, even though we mix original and synthetic images separately. Likewise, Jensen–Shannon Divergence (JSD) loss [Hendrycks2020] fails to significantly boost corruption robustness despite training with 3 times the images. Thus, our method does not combine as broadly with other augmentations as purely rule‑based methods.
From Table 5 it can be seen that our method achieves robustness gains compared to rule-based combinations across various convolutional model architectures. Furthermore, in an ImageNet-to-TIN transfer task (finetuning for 20 epochs with AdamW), our approach boosts a ViT‑B‑16 vision transformers (ViT) robustness while maintaining accuracy.
Table 5: SYN+NST+TA (Ours) outperforms rule-based TA+RE+N augmentation on ResNeXt (RNX), DenseNet (DN) and ImageNet-preactivated Vision Transformer (ViT).
@@ 6 DISCUSSION
Hyperparameter tuning on validation data while using synthetic training data induces a bias. The models generating the synthetic data have been trained to mimic the feature distribution of the training data [karras2022elucidating]. Therefore, the synthetic image features are similar to those of the validation data that is split from the training set. Hence, validation accuracy is artificially inflated and the parameter selection is mislead by overestimating the optimal $\lambda$ for TA+NST. This bias emphasizes the need for separate validation data for benchmark datasets such as CIFAR.
Our methods multiple hyperparameters pose a practical challenge, as standard grid search tuning is likely to miss the varying optima for each dataset. PBT was therefore proposed as an automated tuning method. However, the results were only competitive with respect to accuracy, while the non-constant parameter schedule appeared to hurt the robustness. Different PBT parameters, or automatic tuning schemes that focus on individual data samples may be needed to get the most out of our method.
@@
Table 7: Compute time of one epoch of WRN-28-4 training on a NVIDIA RTX 6000 Ada. Both NST and NoisyMix induce overhead (italics) through GPU-based stylization and JSD loss with 3x images, respectively.
@@ 7.2 Efficient NST during training
Computing in dataloader: Typically, pytorch datasets call single images for transformation, which is computational inefficient for NST. Hence, we implemented a custom dataset and batch sampler class. The dataset instance stylizes and caches the entire next batch whenever its first image is called. This is a less compute efficient, but more memory efficient approach for larger datasets.
Further alternatives include applying NST after data loading, which looses the ability carry out subsequent augmentations in parallel on CPU. Precomputing stylized images and saving them to disk every epoch could save memory. Multiple GPUs would allow parallelization of NST and training. Furthermore, pytorch dataloaders feature a ”collate_fn” function to transform batches of data, however we could not make this work for NST while using multiple CPU workers.

### v3

@@ Abstract
This paper proposes a training data augmentation pipeline that combines synthetic image data with neural style transfer in order to address the vulnerability of deep vision models to common corruptions. We show that although applying style transfer on synthetic images degrades their quality with respect to the common Frechet Inception Distance (FID) metric, these images are surprisingly beneficial for model training. We conduct a systematic empirical analysis of the effects of both augmentations and their key hyperparameters on the performance of image classifiers. Our results demonstrate that stylization and synthetic data complement each other well and can be combined with popular rule-based data augmentation techniques such as TrivialAugment, while not working with others. Our method achieves state-of-the-art corruption robustness on several small-scale image classification benchmarks, reaching 93.54%, 74.9% and 50.86% robust accuracy on CIFAR-10-C, CIFAR-100-C and TinyImageNet-C, respectively.
@@ 1 INTRODUCTION
In recent years, computer vision models have surpassed human performance in tasks such as image classification. However, unlike the human vision system, such models tend to be unstable when faced with small changes to their input data. This is true for both adversarially optimized perturbations [goodfellow2014] and randomly distributed real-world image corruptions [Hendrycks2019a]. Achieving robustness to such changes is an essential technical property of vision models and even a regulatory prerequisite for safety-critical applications [EU_AI_Act_2024].
@@
This study examines the combination of two model-guided data augmentation approaches, synthetic images and Neural Style Transfer (NST), as illustrated in Figure 1. Synthetic images from generative models still exhibit an appearance gap to original images, limiting their effective use in model training [wang2024improving]. We hypothesize that NST can circumvent this gap by exchanging potentially flawed synthetic image textures, making a model learn corruption-robust features from high-quality synthetic object shapes instead.
@@
    We show that NST and synthetic images combine well: Although NST degrades the quality of synthetic images according to the popular Frechet Inception Distance (FID) metric, stylizing synthetic images improves the model performance and is empirically more effective than stylizing original images.
@@ 2 BACKGROUND
where $x,y$ are sampled from a data distribution $D$ and $C$ is a set of corruptions that alter $x$. $C$ can contain real-world [Hendrycks2019a], particularly challenging [Mintun2021], or theoretically motivated image corruptions [wang2021statistically, Siedel2022]. A straightforward way to measure corruption robustness in (1), is to average test accuracies over all corruption sets.
@@
Frechet Inception Distance (FID): Let $X$ be a random vector in a finite-dimensional vector space. We denote its mean as $\mu_{X}=\mathbb{E}[X]$ and its variance as $C_{X}=\mathbb{E}[(X-\mu_{X})(X-\mu_{X})^{\top}]$. Given two random vectors $X$ and $X^{\prime}$, the FID given by
@@ 3.1 Rule-based data augmentation
Mixup [Zhang2018mixup] and Cutmix [Yun2019cutmix] interpolate between two images and their respective labels for training. Mixup boosts model robustness by enforcing a smooth, linear behavior. Noisy Feature Mixup extends Mixup and Noise Injections into the feature space, improving their positive effect. NFM is only one example where combining rule-based methods yields superior results. PyTorch’s top image classification checkpoints use TrivialAugment, Cutmix, Mixup, and Random Erasing [Vryniotis2021]. NoisyMix [Erichson2024] is the top ranking method for achieving corruption robustness for the CIFAR-10/100 datasets on Robustbench [croce2020b] and combines NFM with AugMix and the Jensen–Shannon Divergence (JSD) loss that enforces consistency between augmented samples.
@@ 3.2.1 Synthetic Data Augmentation
Synthetic training data have been applied for both large-scale training [he2023] and in data-scarce fields such as medicine [Frid2018] to boost model performance. Generative models can therefore assist training on downstream tasks such as image classification [azizi2023]. For classifier training, class-conditional image generation, e.g. with [dhariwal2021] or without [ho2021] guidance of a classifier or via text prompts [ramesh2022] enables labeled data generation, thereby eliminating the need for pseudo-labeling [gowal2020uncovering].
Generative adversarial networks have long been used for synthetic data augmentation [zheng2017unlabeled]. Later, diffusion models [ho2020denoising] surpassed generative adversarial networks by delivering superior image quality [rombach2022, karras2022elucidating], typically measured through a lower FID score (see section 2). However, there still exists a significant domain gap between synthetic and original data due to a discrepancy in quality and appearance of synthetic images [wang2024improving]. This gap limits the effectiveness of synthetic data for training.
Several studies focus on improving the adversarial robustness of classifiers in particular through the addition of synthetic data during training [gowal2021improving, rebuffi2021fixing, wang2023better]. In fact, all top adversarially robust models on the popular Robustbench leaderboard [croce2020b] leverage generated data. Although some works hint at benefits for domain- and out-of-distribution generalization [Sandfort2019, wu2023datasetdm], the effect of additional generated training data on corruption robustness as defined in section 2 has to the best of our knowledge not been studied extensively. The top corruption-robust models on Robustbench rely solely on rule-based data augmentation [croce2020b].
@@ 3.2.2 Neural Style Transfer
NST was used as a data augmentation method by [geirhos2018imagenettrained] when introducing a stylized version of the ImageNet dataset. Training with this dataset improves corruption robustness by forcing a convolutional neural network (CNN) to recognize object shapes instead of overfitting on textures, as those are random after stylization. However, [geirhos2018imagenettrained] trained on a fixed stylized version of the dataset. Ideally, we want an adaptive data augmentation scheme that allows a new stylization view every training epoch. [jackson2019style] did such data augmentation using the NST approach in [ghiasi2017exploring]. However, they did not evaluate corruption robustness. Neither of the two works has used their methods to fix the appearance gap of synthetic training data by exchanging unrealistic textures.
@@ 4 METHODOLOGY
Our data augmentation approach is to the best of our knowledge the first to explore a combination of synthetic image data augmentation and NST for maximizing the corruption robustness of image classifiers. To this end, our approach makes use of original, synthetic and style images as can be seen from Figure 1.
@@ 4.2 Stylization
  • $\lambda_{o}$ and $\lambda_{s}$ are independent probabilities that a synthetic image or an original image is stylized, respectively (see Figure 1).
@@
Table 1: FID of synthetic images over stylization ratio $\lambda_{s}$ for 50.000 synthetic images (100.000 for TinyImageNet). $\alpha_{s}$ is random uniformly drawn from [0.1,1.0] as is optimal in our experiments. Original denotes the FID before any pre-selection against the original reference.
@@ 5.2 Stylization ruins FID but improves model training
Let $D$ be the marginal feature distribution of the original training data. The mean and variance of the features $X$ of the original training data are given by $\mu=\mathbb{E}_{X\sim D}[X]$ and $C=\mathbb{E}_{X\sim D}[(X-\mu)(X-\mu)^{\top}]$. The mean and variance of the feature distribution $D^{\prime}$ of the synthetic dataset are given by $\mu^{\prime}$ and $C^{\prime}$. Let $D^{\prime}_{s}$ be a joint distribution of style images and synthetic data.
@@
Figure 2: Test accuracy, robustness, and FID of stylized synthetic images plotted over varying synthetic stylization ratios $\lambda_{s}$. The accuracy and robustness maxima are not aligned with the minimum FID of the synthetic training data; hence, FID is not a good predictor for the usefulness of synthetic images for data augmentation.
@@ 5.3 Parameter Tuning
Parameter $\alpha$: Figure 3 plots validation‑set tuning of $\alpha_{o}$ and $\alpha_{s}$ for TinyImageNet. For all our datasets, contrary to results in [jackson2019style] where $\alpha=0.5$ is optimal, we achieve best accuracy and robustness on original data at $\alpha_{o}=1.0$. On synthetic data, $\alpha_{s}$ is in a trade‑off, where increasing it too far boosts robustness but reduces accuracy. We choose to select $\alpha_{s}$ randomly drawn from $[0.1,1.0]$ as a balanced choice.
@@
Figure 5: Robustness and accuracy maps depending on the ratio of stylization on original ($\lambda_{o}$, x-axis) and synthetic ($\lambda_{s}$, y-axis) images. Values are averaged across 5 runs. The synthetic image ratio $\lambda$ is 0.5. The points mark the lambda combination that is optimal for that dataset with respect to the mean of accuracy and robustness. On CIFAR in general, but also for optimal robustness on TIN, $\lambda_{o}<\lambda_{s}$.
@@
Population-based training:
@@
In addition to the fragmentary grid search for individual hyperparameters, we generate an optimized augmentation schedule using population-based training (PBT) [jaderberg2017population, ho2019population]. PBT trains multiple trials in parallel, promotes successful over unsuccessful checkpoints every predefined interval and perturbs the augmentation hyperparameters. It then tracks the most effective schedule over the training epochs. Figure 7 shows the resulting schedules for the lambda parameters, while the alpha parameters are set to constant values as derived from prior tuning. We apply PBT on train-validation data without resampling of parameters, with a perturbation interval of 1 epoch, after which trials are ranked according to the mean of accuracy and robustness, and 24 overall trials. We then replayed the schedule on train-test data. The schedule confirms for the CIFAR datasets that in tendency, synthetic data can take more stylization than original data, hence $\lambda_{s}>\lambda_{o}$. Towards the low learning rate sections of the training schedule at 140 and 300 epochs, there is a tendency towards more policy changes and higher stylization probabilities. The best synthetic ratio $\lambda$ clearly trends down from the start value of 0.5 that was found optimal before.
Table 2: Ours (SYN+NST+TA) outperforms other data augmentation methods. *NoisyMix is a 300 epoch rerun of the original implementation without early stopping, IPMix results are from the original paper. Both feature different parameters, IPMix uses the deeper WRN-40-4 architecture [Erichson2024, huang2024ipmix].
@@ 5.4 Ablation and Comparison Studies
Table 4 shows further ablations, adding popular rule‑based augmentations to the SYN+NST+TA baseline. Adding Random Erasing and noise in input and feature space in most cases further improves robustness. Prior work demonstrated that the listed additional approaches synergize well in a purely rule‑based setting [Vryniotis2021, Erichson2024]. However, combined with our approach, Mixup and CutMix degrade performance, even though we mix original and synthetic images separately. Likewise, JSD loss as in [Hendrycks2020] fails to significantly boost corruption robustness despite training with 3 times the images. Thus, our method does not combine as broadly with other augmentations as purely rule‑based methods.
From Table 5 it can be seen that our method achieves robustness gains compared to rule-based combinations across various convolutional model architectures. Furthermore, in an ImageNet-to-TIN transfer task (finetuning for 20 epochs with AdamW), our approach boosts ViTs robustness while maintaining its accuracy.
Table 5: SYN+NST+TA (Ours) outperforms rule-based TA+RE+N augmentation on ResNeXt (RNX), DenseNet (DN) and ImageNet-preactivated ViT.
@@ 6 DISCUSSION
Hyperparameter tuning on validation data while using synthetic training data induces a bias. The models generating the synthetic data have been trained to mimic the feature distribution of the training data [karras2022elucidating]. Therefore, the synthetic image features are similar to those of the validation data split from the training set. Hence, validation accuracy is artificially inflated and the parameter selection is misled by overestimating the optimal $\lambda$ for TA+NST. This bias emphasizes the need for separate validation data for benchmark datasets such as CIFAR.
Our method’s multiple hyperparameters pose a practical challenge, as standard grid search tuning is likely to miss the varying optima for each dataset. PBT was therefore proposed as an automated tuning method. However, the results were only competitive with respect to accuracy, while the non-constant parameter schedule appeared to hurt the robustness. Different PBT parameters, or automatic tuning schemes that focus on individual data samples may be needed to get the most out of our method.
@@
Table 7: Compute time of one epoch of WRN-28-4 training on a NVIDIA RTX 6000 Ada. Both NST and NoisyMix induce overhead through GPU-based stylization and JSD loss with 3x images, respectively.
@@ 7.2 Efficient NST during training
Computing in dataloader: Typically, PyTorch datasets call single images for transformation, which is computationally inefficient for NST. Hence, we implemented a custom dataset and batch sampler class. The dataset instance stylizes and caches the entire next batch whenever its first image is called. This is a less compute efficient, but more memory efficient approach for larger datasets.
Further alternatives include applying NST after data loading, which loses the ability carry out subsequent augmentations in parallel on CPU. Precomputing stylized images and saving them to disk every epoch could save memory. Multiple GPUs would allow parallelization of NST and training. Furthermore, PyTorch dataloaders feature a ”collate_fn” function to transform batches of data, however we could not make this work for NST while using multiple CPU workers.